Question  2

Find the derivative of the function Y= x3 + x2 /x


a. x2/x + x/x


b. x2+ x /x


c. 2x + 1


d. 3x4 + 2x3



1
Expert's answer
2021-06-23T13:33:23-0400

"y = \dfrac{x³ + x²}{x}=
\\ u = x³ + x²\\
du/dx = 3x² + 2x\\
v = x
\\dv/dx = 1

\\
\dfrac{dy}{dy} = \dfrac{Vdu/dx-Udv/dx}{V²}"


"\dfrac{dy}{dx} =\dfrac{x(3x²+2x)-(x³+x²)(1)}{x²}"


"= \dfrac{3x³+2x²-x³-x²}{x²} =\dfrac{2x³+x²}{x²}"


"=2x+1"


"a.\ \ y = x +1\\
\quad \dfrac{dy}{dx} = 1(x⁰) = 1"


"b.\ \ y = 2x +1\\
\quad \dfrac{dy}{dx} = 2(x⁰) +0 = 2"


"c.\ \ y = 2x +1\\
\quad \dfrac{dy}{dx} = 2(x⁰)+ 0 = 2"


"d.\ \ y = 3x⁴ +2x³\\
\quad \dfrac{dy}{dx} = 3(4x³) + 6(x²) = 12x³ + 6x²"